A shipment of nine printers contains two they are defective. Find the probability that a sample of size 2, drawn from nine, will not contain a defective printer


1
Expert's answer
2022-03-09T14:28:01-0500

The shipment contains 9 printers and 2 are defective. 2 printers are chosen at random.


2 out of 9 can be chosen in "n=C_9^2" ways (Total number of outcomes)


Number of favourable outcomes (not containing defective)


"m=C_{9-2}^2=C_7^2"


The Probability that none is defective is


"p =\dfrac{m}{n}=\dfrac{C_7^2}{C_9^2}=\dfrac{7}{12}"


"=\dfrac{7}{12}"